OpenStudy (anonymous):

i'm trying to determine if n^1/2 logn^1/2 is O(n). where do i start?

OpenStudy (neer2890):

use basic Big O notation: f(n) = O g(n) if and only if \[0\le f(n)\le c1 g(n) \] where c1 is an arbitrary constant.

OpenStudy (neer2890):

here, \[f(n)=n ^{\frac{ 1 }{ 2 }}\log n ^{\frac{ 1 }{ 2 }}\] and g(n)=O(n). now find some constant c1 which fits the above equation if it exists then you can say that f(n)=O g(n) otherwise not.